# Enhancing Consistency-Based Image Generation via Adversarialy-Trained Classification and Energy-Based Discrimination

### v1

@@ 3.2 Sampling
Algorithm 2 Boosting Images via Joint Classifier-Discriminator
@@ 4 Experiments
In this section, we present an evaluation of the proposed boosting method. We begin by discussing the results obtained by our method compared to BIGROC [6]. Subsequently, we conduct an ablation study to further analyze various components of the model, as described in Section  3.
@@ 4.1 Experimental Settings
In the training phase, we use the pre-trained robust RN50 model [21] and replace its fully connected head with a randomly initialized one. We configure the PGD parameters for the attack on fake images with an $\varepsilon$ of 10.0 and a step size of 0.1, incrementing the number of steps every 1000 iterations. For real images, the PGD attack is set up with an $\varepsilon$ of 1.0, a step size of 0.25, and limited to 4 steps. The model is trained on 4 NVIDIA GeForce RTX 3090 GPUs. In the sampling process, we apply the loss function outlined in Equation  13, the step size is set as $\varepsilon/T$, where $T$ is the number of steps. The amount of steps is adjusted according to the generative model we enhance; further technical details can be found in Appendix  D.
Given the synthesized images obtained from consistency models, $X_{\theta}$, we apply our method to generate enhanced samples, $X_{boosted}$, as explained in Algorithm  2. We evaluate our method on the two consistency methods: CD and CT [26]. In all the experiments, we use the same joint model acquired from the training process as explained in  3.1. In each experiment, we initialize $\varepsilon$ as mentioned above and adjust only the number of steps taken in the PGD process, which defines the allowable distance to move from the original sample.
@@ 4.2 Experimental Results
We analyze the perceptual-quality boosting performance using Fréchet Inception Distance ( [12], lower is better), and Inception Score ( [23], higher is better). Table 1 summarizes our results on ImageNet 64x64  [3]. First, we provide the original FID and IS results as described in the original Consistency Models paper [26]. Then, we apply the boosting algorithm with a robust classifier and our joint model. As can be seen, the obtained results indicate that the proposed approach is more beneficial compared to a refinement via a robust classifier alone (i.e. BIGROC  [6]). Observe the critical effect of the robustness in the training of our joint model, which leads to much improved results. This boost in quality is obtained across both CT and CD and when operated using either one or two generative steps.
In terms of computational complexity, applying one PGD step takes only 0.02 seconds, while each inference step of Consistency (CT or CD) takes 0.55 seconds. Therefore, 25 PGD steps of our algorithm are roughly equivalent to a Consistency inference. This suggests that our approach could be interpreted as providing a continuum tradeoff between complexity and perceptual quality.
For reference, the work reported in [17] addressed the same task of boosting CT synthesis. They approached their task by wrapping the Consistency with a GAN structure, in which the CT network itself serves as a generator to be retrained, being governed by a discriminator via adversarial training. Note that their training is much more involved than ours, and much heavier in computational complexity, due to the size of the CT network and the adversarial loss being used. [17] obtained an FID of $10.6$, which is slightly better than BIGROC, but falls short when compared with the boosting of our joint machine.
Table 1: FID and IS results of Consistency and the results after boosting the synthesized images via a robust classifier, and with a robust or vanilla joint model, as proposed in this work.
@@
Boosting Method Inference Steps FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ )
Consistency Training 1 13.00 28.83
Robust Classifier 10.71 36.38
Non-Robust Joint Model 12.30 35.62
Robust Joint-Model 9.60 50.80
Consistency Training 2 11.12 27.28
Robust Classifier 9.65 33.55
Non-Robust Joint Model 11.28 33.16
Robust Joint-Model 8.51 39.40
Consistency Distillation 1 6.20 39.87
Robust Classifier 5.55 44.65
Non-Robust Joint Model 6.33 45.55
Robust Joint-Model 4.95 51.98
Consistency Distillation 2 4.69 42.28
Robust Classifier 4.20 46.78
Non-Robust Joint Model 4.66 46.37
Robust Joint-Model 3.84 51.12
Table 2 presents similar results for three additional generative models, in which we operate and boost their output images using the joint model as described above (no retraining is employed). The obtained results imply that the joint model has a generalization power and achieves good improvements even though it was trained on Consistency-based images only. Future work is required to assess adaptation of our classifier-discriminator model for better boosting these generative methods.
Table 2: FID results for three generative methods and the results after boosting the synthesized images via a robust classifier and the joint model proposed.
Boosting Method FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ )
Admnet  [4] 2.61 46.78
Robust Classifier 2.40 50.13
Robust Joint-Model 2.31 53.00
BigGAN  [2] 4.06 44.94
Robust Classifier 3.92 48.76
Robust Joint-Model 3.83 51.75
Iddpm  [22] 2.92 45.62
Robust Classifier 2.67 48.91
Robust Joint-Model 2.55 51.76
@@ 4.3 Ablation Study
As explained in Section  3.2, the choice of the loss function in Algorithm  2 is crucial. The results in Table  3 illustrate that utilizing only Cross-Entropy loss produces results that are similar to those achieved using BIGROC  [6], as expected. The joint model demonstrates substantial discriminative power, which further enhances FID scores when Binary Cross-Entropy loss is applied. In exploring the integration of these two components, we experimented with the loss function proposed in equation  12. Our findings indicate that this approach improves upon Cross-Entropy loss, though it does not consistently surpass Binary Cross-Entropy loss. Consequently, we recommend adopting the loss outlined in Equation  13, which shows an additional improvement in FID by approximately 0.1.
Table 3: FID results with varying loss functions as described in section  3.2.
@@ 5 Conclusion
Limitations: In this work we have used the RN50 architecture for the joint model, as commonly employed in robust classification tasks. We note that this architecture is relatively simple and small, compared to the more complex ones proposed in recent years. As a consequence, our study is unavoidably constrained by the capabilities of the chosen RN50 model. Another limitation in our work is the reliance of our training on Consistency generated images only. Expanding the training-set to include images from additional generative models could further improve the overall boosting effect that we document in this paper.
@@ Appendix D Supplementary Results
In Figure 1 we present qualitative results demonstrating that the “boosted”results indeed look better perceptually. The enhancement is visible on the edges and textures and leads to sharper and more high-contrast images, creating a more perceptually pleasing outcome.
Figure 1: Qualitative demonstration of the boosting algorithm results. Top: Consistency Training images. Bottom: boosted images after applying our proposed method.
@@
For reproducibility, we elaborate on the amount of steps and the step size used for boosting each generative model in Table  4.

### v2

@@ 3.2 Sampling
Algorithm 2 Boosting Images via PGD Guided by a Joint Classifier-Discriminator
@@
An alternative approach is to sample using Stochastic Gradient Langevin Dynamics, similarly to  grathwohl2019your , as described in Algorithm  3. To further investigate the properties of SGLD, we explored replacing PGD with SGLD during both inference and training. During training, we found that optimizing with SGLD lacks the robustness achieved with PGD and suffers from convergence difficulties. However, during inference, SGLD surpasses PGD, as shown in Table  3. The benefit of using SGLD during inference can be attributed to its ability to better explore multiple modes compared to methods like PGD due to it’s stochastic nature.
Algorithm 3 Boosting Images via SGLD Guided by a Joint Classifier-Discriminator
Input:
  • Robust Joint Classifier-Discriminator $f_{\theta}(x)$,
  • Input image $x$ of a target class $c$, and
  • Algorithm’s parameters: Step-size $\alpha$, No. of iterations $T$, noise level $\sigma$.
for $t$ from $1$ to $T$ do
     $x_{t}=x_{t-1}-\alpha\cdot\nabla_{x_{t-1}}f_{\theta}(x_{t-1})\left[c\right]+\sigma\cdot\mathcal{N}(0,\textbf{I})$.
end for
Output: $x_{T}$.
@@ 4 Experiments
In this section, we present an evaluation of the proposed boosting method. We begin by discussing the results obtained by our method compared to BIGROC ganz2022bigroc . Subsequently, we conduct an ablation study to further analyze various components of the model, as described in Section  3.
@@ 4.1 Experimental Settings
During training, we use the pre-trained robust ResNet-50 (RN50) model madry2018towards as a backbone, with a randomly initialized classification head. We configure the parameters of the Projected Gradient Descent (PGD) attack for fake images with an $\varepsilon$ of 10.0 and a step size of 0.1, increasing the number of PGD steps every 1,000 training iterations to enhance adversarial robustness over time. For real images, the PGD attack is set with an $\varepsilon$ of 1.0, a step size of 0.25 and 4 steps. The model is trained on four NVIDIA GeForce RTX 3090 GPUs. The same configuration applies to the training of the Wide-ResNet-50-2 (Wide-RN) model. In the sampling process, we apply the loss function outlined in Equation  14 and the step size is set as $0.1$. The amount of steps is adjusted according to the generative model we enhance; further technical details can be found in Appendix  D.
@@ 4.2 Experimental Results
We analyze the perceptual-quality boosting performance using Fréchet Inception Distance  heusel2017gans , Inception Score  salimans2016improved , Precision and Recall. Table 1 summarizes our results on the ImageNet 64×64 dataset deng2009imagenet . Initially, we present the original FID and IS results as reported in the Consistency Models paper song2023consistency . We then apply our boosting algorithm using both a robust classifier and our joint model. The results demonstrate that our proposed approach offers greater benefits compared to refining with a robust classifier alone (i.e., BIGROC ganz2022bigroc ). Notably, the proposed training of our joint model leads to significantly improved results. This boost in image quality is observed across both Consistency Training (CT) and Consistency Distillation (CD) methods, and when using either one or two generative steps. In terms of efficiency, one PGD step requires only 0.02 seconds, while a single Consistency (CT or CD) inference step takes 0.55 seconds, making 25 PGD steps comparable in cost to a Consistency inference. This suggests that our approach could be interpreted as providing a continuum trade-off between complexity and perceptual quality. For comparison,  kong2023act used a GAN-based approach to improve CT synthesis with adversarial training, achieving an FID of 10.6, which slightly outperforms BIGROC but falls short of our joint model’s performance boost. Note that their training is much more extensive than ours, and much heavier in computational complexity, due to the size of the CT network and the adversarial loss being used.
Table 1: Quantitative comparison of image generation performance on ImageNet 64×64 using Consistency Models and our proposed boosting methods with robust classifiers and joint models.
@@
Boosting Method Inference Steps FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
Consistency Training 1 13.00 28.83 0.71 0.41
RN50 Robust Classifier 10.71 36.38 0.70 0.47
Wide-RN Non-Robust Joint 12.30 35.62 0.64 0.46
RN50 Robust Joint 9.60 50.80 0.76 0.39
Wide-RN Robust Joint 8.83 55.67 0.73 0.45
Consistency Training 2 11.12 27.28 0.68 0.54
RN50 Robust Classifier 9.65 33.55 0.68 0.54
Wide-RN Non-Robust Joint 11.28 33.16 0.63 0.54
RN50 Robust Joint 8.51 39.40 0.74 0.49
Wide-RN Robust Joint 7.98 46.33 0.71 0.52
Consistency Distillation 1 6.20 39.87 0.67 0.63
RN50 Robust Classifier 5.55 44.65 0.68 0.61
Wide-RN Non-Robust Joint 6.33 45.55 0.64 0.60
RN50 Robust Joint 4.95 51.98 0.72 0.58
Wide-RN Robust Joint 4.84 58.73 0.70 0.60
Consistency Distillation 2 4.69 42.28 0.68 0.63
RN50 Robust Classifier 4.20 46.78 0.69 0.63
Wide-RN Non-Robust Joint 4.66 46.37 0.68 0.63
RN50 Robust Joint 3.84 51.12 0.72 0.60
Wide-RN Robust Joint 3.78 55.13 0.71 0.61
@@ 4.3 Ablation Study
As explained in Section  3.2, the choice of the loss function in Algorithm  2 is crucial. The results in Table  2 illustrate that utilizing only Cross-Entropy loss produces results that are similar to those achieved using BIGROC  ganz2022bigroc , as expected. The joint model demonstrates substantial discriminative power, which further enhances FID scores when Binary Cross-Entropy loss is applied. In exploring the integration of these two components, we experimented with the loss function proposed in equation  13. Our findings indicate that this approach improves upon Cross-Entropy loss, though it does not consistently surpass Binary Cross-Entropy loss. Consequently, we recommend adopting the loss outlined in Equation  14, which shows an additional improvement in FID by approximately 0.1.
Table 2: FID and IS results with varying loss functions as described in section  3.2.
@@
We also compare sampling methods — Projected Gradient Descent (PGD) and Stochastic Gradient Langevin Dynamics (SGLD) — as discussed in Section 3.2 and presented in Table 3. Our experiments reveal that SGLD offers advantages over PGD in certain contexts, particularly in terms of sampling diversity and fidelity.
Table 3: Quantitative comparison of our proposed boosting methods using Wide-ResNet with PGD and SGLD sampling techniques applied to Consistency Models.
Boosting Method Inference Steps FID ()($\downarrow$)( ↓ ) IS()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
Consistency Training 1 13.00 28.83 0.71 0.41
Wide-RN$+$PGD 8.83 55.67 0.73 0.45
Wide-RN$+$SGLD 8.71 59.76 0.73 0.46
Consistency Training 2 11.12 27.28 0.68 0.54
Wide-RN$+$PGD 7.98 46.33 0.71 0.52
Wide-RN$+$SGLD 7.64 46.83 0.71 0.53
Consistency Distillation 1 6.20 39.87 0.67 0.63
Wide-RN$+$PGD 4.84 58.73 0.70 0.60
Wide-RN$+$SGLD 4.65 59.00 0.71 0.60
Consistency Distillation 2 4.69 42.28 0.68 0.63
Wide-RN$+$PGD 3.78 55.13 0.71 0.61
Wide-RN$+$SGLD 3.58 56.46 0.72 0.61
To assess the generalization capability of our joint model, we evaluated it on images generated by BigGAN and ADM-G at resolutions of 128×128 and 256×256 pixels (see Table 4). Despite being trained exclusively on images from Consistency Models, our joint model demonstrates strong generalization across different generative models and resolutions. This suggests that our model can enhance images from a variety of generative sources without requiring retraining or fine-tuning for each specific case.
Table 4: Quantitative comparison for various generative models at different resolutions (64×64, 128×128, and 256×256), before and after applying our proposed boosting method using Wide-ResNet with SGLD.
Boosting Method FID ()($\downarrow$)( ↓ ) IS ()($\uparrow$)( ↑ ) Precision ()($\uparrow$)( ↑ ) Recall ()($\uparrow$)( ↑ )
Adm 64x64 2.61 46.78 0.73 0.63
Wide-RN$+$SGLD 2.08 58.41 0.75 0.60
BigGAN 64x64 4.06 44.94 0.79 0.48
Wide-RN$+$SGLD 3.67 62.79 0.81 0.45
Iddpm 64x64 2.92 45.62 0.73 0.62
Wide-RN$+$SGLD 2.25 62.26 0.76 0.59
Adm-G 128x128 2.97 141.47 0.78 0.59
Wide-RN$+$SGLD 2.34 271.95 0.80 0.58
BigGAN 128x128 6.02 145.83 0.86 0.34
Wide-RN$+$SGLD 5.68 236.88 0.86 0.34
Adm-G 256x256 4.58 186.83 0.81 0.52
Wide-RN$+$SGLD 3.17 301.06 0.83 0.53
BigGAN 256x256 7.03 202.64 0.87 0.27
Wide-RN$+$SGLD 6.16 275.25 0.88 0.28
@@ 5 Conclusion
Limitations: In this work we have used the RN50 and Wide-RN 50-2 architectures for the joint model, as commonly employed in robust classification tasks. We note that this architectures are relatively simple and small, compared to the more complex ones proposed in recent years. As a consequence, our study is unavoidably constrained by the capabilities of the chosen models. Another limitation in our work is the reliance of our training on Consistency generated images only. Expanding the training-set to include images from additional generative models could further improve the overall boosting effect that we document in this paper.
@@ Appendix D Supplementary Results
In the following figures we present qualitative results demonstrating that the “boosted” results indeed look better perceptually. The enhancement is visible on the edges and textures and leads to sharper and more high-contrast images, creating a more perceptually pleasing outcome. Note that we recommend zooming in the images to see the changes clearly.
For reproducibility, we elaborate on the amount of steps and the step size used for boosting each generative model in Table  5.
Figure 1: Top: Images generated by Consistency Models trained on ImageNet 64x64. Bottom: Refined images by applying our algorithm.
Figure 2: Top: Refined images by applying our algorithm. Bottom: Images generated by Guided Diffusion trained on ImageNet 256x256.
Figure 3: Top: Images generated by Guided Diffusion trained on ImageNet 256x256. Bottom: Refined images by applying our algorithm.
Figure 4: Top: Images generated by Guided Diffusion trained on ImageNet 128x128. Bottom: Refined images by applying our algorithm.